OpenStudy (anonymous):

If a skier skis down the hill at 70.0 km/h for 20.0 s, and then skis up a slope at 15.0 km/h for 10.0 s, what is the skier's average velocity? (Remember that the question asks for velocity, not speed.)

OpenStudy (abmon98):

Average Velocity=Total Distance Traveled/Total Time Taken V=D/T D=VT (70*20)+(15*10)/20+10